OpenStudy (anonymous):

I have this question for MultiA: f(x,y,z)=sin(xy)cos(z)+xyz, a)FInd the rate of change of f at point P=(pi,0,pi) in the direction of the vector v=i +j - k b) In which direction does f increase more rapidly at the point P c) What is the max rate of change of f at P d)Determine the equation of the tangent plane to the level surface of f at P?

OpenStudy (anonymous):

The directional derivative at P_0 is given by: \[\vec{\nabla}f(x,y,z)|_{P_0} \cdot \vec{u}_{\vec{v}}\] Where u_v denotes a unit vector in the v direction. So first: \[| \vec{v}|=\sqrt{1^2 + 1^2 + 1^2}=\sqrt{3} \implies \vec{u}_{\vec{v}}=\frac{1}{\sqrt{3}}<1,1,-1>\] Then: \[\vec{\nabla}f(x,y,z)=<\frac{\partial f}{\partial x}, \frac{\partial f}{\partial y}, \frac{\partial f}{\partial z}>=<y \cos(xy)\cos(z)+yz,\] \[x \cos(xy)\cos(z) + xz, -\sin(xy) \sin(z) + xy>\] Evaluate the gradient at P. Form the dot product and you're done :P The direction is of most increase is the gradient. Most rapid increase is the magnitude of the gradient: \[||\vec{\nabla}f(x,y,z)|_P||\] And to find the plane you have the perpendicular vector (most increase) then: \[\hat{n} \cdot (x-P_1,y-P_2,z-P_3)=0; P=(P_1,P_2,P_3)\] Using normal tangent vector definition.